A flat surface with area 0.14m^2 lies in the x-y plane, in a uniform electric field given by E= 5.1i^+2.1j^+3.5k^ kN/C. find the flux through this surface
Answer #1

Flux = ∫E dA = E·A = EA cos(θ)

Since the surface lies in the xy-plane, it's area vector is pointed in the z-direction, so only the z-component of the electric field contributes to the flux.

Flux = E·A = (5.1, 2.1, 3.5)kN/C · (0, 0, 0.14) m2 = 490. N m2 / C